Part A Three 20-cm-long rods form an equilateral triangle. Two of the rods are charged to 12 nC, the third to -12 n What is the electric field strength at the center of the triangle? Express your answer to two significant figures and include the appropriate units. EnetVaue Units Submit My Answers Give Up
Answer #1

r = sqrt(3)/6*20 cm =0.0577 m
tan(30) = r/(L/2)
E' = kQ/{r*sqrt[(L/2)^2 + r^2]} = 8.99e9 * 12e-9 / {0.0577 * sqrt[(0.20/2)^2 + 0.003]} = 16398 N/C (this is the electric field strength for a single rod at the center)

|E'| = 16398 N/C
E = 2|E'|sin(30) + |E'| = 32796 N/C